### v1

@@ 1 Introduction
There has been a long history of textures proposed for the quark mass/Yukawa matrices (MMs), both as stand-alone ansatze, and/or as the result of (or motivation for) models which can constrain them. The traditional guiding principle for such textures has been to reduce the number of free parameters (a priori 36) to a number smaller than the ten observables, in order to find predictive constraints among the measurable quantities, especially relationships between the quark masses and mixings fritzsch1 fritzsch2 fritzsch3 texZero , which each manifest marked hierarchical structure. Methods have included using symmetries to enforce texture-zeroes zeroesSymms , and/or mechanisms to produce a hierarchy among the matrix elements fnMech . Recently, ideas have turned to the realm of modular forms modular0 modular1 , where again approaches have been found to enforce zeroes modularZeroes and/or hierarchies modularHierarchies . We introduce here a new MM texture based purely on the CKM and quark mass phenomenology, which is modestly predictive, fits the data, and has some novel and potentially interesting features.
The closeness of the (beauty-down) Unitarity Triangle (UT) UT1 UT2 to a right-angled triangle is striking rightUT AKMS , and it is also notable that its three internal angles pdg , [^1] We use the measured value for $\beta$ and the CKM fit values pdg for $\alpha$ and $\gamma$, since these give the smallest uncertainties.
@@ 2 The texture
The proposed new texture is common to both the up and down quark MMs, constraining them relative to those of the SM. Without loss of generality, we take the MMs to be Hermitian. We adopt the popular texture zero in the $13$ and $31$ elements fritzsch1 fritzsch2 fritzsch3 texZero , so that the smallest mixing elements result from the non-commutation of the (small) diagonalising transformations in the $12$ and $23$ subspaces. In order both to introduce viable hierarchies among the MMs’ elements, and to include $CP$ violation, we adopt two small, complex expansion parameters, [^2] Throughout this paper, where ambiguity can arise, bold font is used to denote complex quantities, and the corresponding light font symbols, their moduli. Complex conjugation is indicated by the conventional symbol. $\bm{\lambda_{u}}$ and $\bm{\lambda_{d}}$, powers of which are inserted into the up and down quark MMs respectively, each playing a role similar to Wolfenstein’s $\lambda$ parameter in the CKM matrix wolfenstein , but distinct for each of the two MMs. Thus, we express both MMs as follows:
@@
with the normalisations $n_{q}\simeq m_{t},m_{b}$ for $q=u,d$ respectively, the real coefficients $(A_{0},b,c_{u},c_{d})$ are of magnitude $\lesssim{\cal O}(1)$, and the ‘
@@
’ symbol means “the complex conjugate of the diagonally-opposite entry”. The equality in the coefficients between the 12 and 22 elements is conventional, any a priori difference having been absorbed into the normalisations of $\bm{\lambda_{q}}$, by virtue of the different powers involved.
The overall normalisations of the $\bm{\lambda_{q}}~{}(q=u,d)$ are free to vary, but we anticipate that their complex sum has magnitude close to the value of Wolfenstein’s expansion parameter:
@@
The constraint, eq. (5), is the most significant innovation in this paper. Specifically, in our texture, $(\bm{\lambda_{u}}/\bm{\lambda_{d}})^{*}$ is identifiable in leading approximation with the complex ratio of the two sloped sides of the UT as we show below. Thus $\arg{(\bm{\lambda_{u}}/\bm{\lambda_{d}})}$ corresponds to the supplement of the UT angle $\alpha$, the constrained value ensuring $\alpha\simeq\tfrac{\pi}{2}$; the ratio $|\bm{\lambda_{u}}/\bm{\lambda_{d}}|$ then corresponds similarly to $\tan{\beta}$ (while simultaneously determining the asymmetry between the mass hierarchies in the two charge sectors). Thus eq. (5) applied in our texture, eq. (3), yields eq. (2). Moreover, the precise values for the ratio asserted by eq. (5) imply and are implied by two novel symmetries of the $M_{q}^{0}$, as we will also show.
The MM texture, eq. (3), uses five free parameters, $A_{0}$, $b$, $c_{u}$, $c_{d}$ and $\lambda_{0}$ to fit the three quark mixing angles, the $CP$ phase, and the four quark mass ratios, making it modestly predictive. A more general parameterisation would allow the $(A_{0},b)$ parameters to be different for $q=u,d$, but we find that it is possible to fit all the quark mass and mixing observables with the parameter-pair $(A_{0},b)$, common to both mass matrices. Thus an $up/down$ (weak isospin) reflection symmetry is respected by the real parameters ($A_{0}$, $b$), and broken only by the $\bm{\lambda_{q}}$, the pair of parameters $c_{q}$ which govern $m_{u}$ and $m_{d}$, and by the matrix normalisations, $n_{q}$ in eq. (3).
@@ 3 Leading-order solution
where $\lambda$, $\rho$, $\eta$ and $A$ are the Wolfenstein parameters wolfenstein (and eq. (13) explains our choice of symbol $A_{0}$ in eq. (3)). This confirms that at leading order, $\bm{\lambda_{u}}^{*}$ and $\bm{\lambda_{d}}^{*}$ are each proportional to one of the “sloped” sides of the UT (since, in the CKM phase convention, $V_{ud}\simeq V_{tb}\simeq 1$). We can thus immediately justify the claims made below eq. (4):
@@
$\Rightarrow\arg\left({\frac{\bm{\lambda_{u}}}{\bm{\lambda_{d}}}}\right)$ $=\arg\left({\frac{V_{td}}{V_{ub}^{*}}}\right)\simeq\pi-\alpha~{}{\rm and}~{}\left|\frac{\bm{\lambda_{u}}}{\bm{\lambda_{d}}}\right|\simeq\tan{\beta},$ (15)
where the last near-equality is a consequence of the near right-angle in the resulting UT (the definition of $\alpha$ is from the PDG pdg ).
Still in the PDG phase convention where $\bm{\lambda_{d}}+\bm{\lambda_{u}}=\lambda_{0}$, it is useful to define the complex constants: [^5] The “bar” in $\bm{\overline{z}_{0}}$ is simply to identify this variable as the one-complement of $\bm{z_{0}}$ and should not be confused with the bar appearing above the “reduced” Wolfenstein parameters $\overline{\rho}$ and $\overline{\eta}$.
@@
The apex of the standard UT is simply given by pdg burasetal :
@@ 4 Numerical fit to the data
For the fit to the data, a full diagonalisation is performed at high numerical precision. The CKM parameters are taken from the latest PDG compilation pdg . Significant recent progress in quark mass determinations using lattice QCD flag has led to dramatic reductions in the uncertainties on the quark mass ratios (which we take as experimental inputs [^7] Lattice quark mass determinations are calibrated to experimentally-determined meson masses.). We follow the approach of bb , updating the inputs with the latest lattice determinations compiled in the PDG review on quark masses pdg . These inputs are initially quoted in the second column of table 1 renormalised to the scale $\mu=m_{t}=173.1$ GeV in the $\overline{\rm MS}$ scheme. A $\chi^{2}$ fit to the data is performed, varying the five free parameters of the MMs to find their preferred values, which are given in the table caption.
@@
$|m_{u}/m_{c}|\,(\times 10^{3})$ $2.00\pm 0.05$ Unchanged $2.00$ 0.000.00~{}~{}0.000.00
$|m_{d}/m_{s}|\,(\times 10^{2})$ $4.97\pm 0.06$ $4.97$ 0.030.03~{}~{}0.030.03
$m_{c}/m_{t}\,(\times 10^{3})$ $3.67\pm 0.04$ $3.46\pm 0.03$ $3.46$ $-0.01$
$m_{s}/m_{b}\,(\times 10^{2})$ $1.854\pm 0.007$ $1.968\pm 0.008$  1.9681.968~{}\,1.9681.968 0.020.02~{}~{}0.020.02
$\lambda$ $0.22501\pm 0.00068$ Unchanged 0.224990.22499~{}~{}~{}~{}0.224990.22499 $-0.03$
$A$ $0.826^{+0.016}_{-0.015}$ $0.877^{+0.017}_{-0.016}$ 0.8760.876~{}~{}0.8760.876 $-0.09$
$\overline{\rho}$ $0.1591\pm 0.0094$ Unchanged 0.15190.1519~{}~{}~{}0.15190.1519 $-0.76$
$\overline{\eta}$ $0.3523^{+0.0073}_{-0.0071}$ 0.34770.3477~{}~{}~{}0.34770.3477 $-0.65$
@@
$\beta\,(^{\circ})$ $22.6\pm 0.4$ 22.322.3~{}22.322.3 $-0.75$
$\gamma\,(^{\circ})$ $65.7\pm 1.3$ 66.466.4~{}66.466.4   0.54
$J_{CP}\,(\times 10^{5})$ $3.12^{+0.13}_{-0.12}$ $3.49^{+0.15}_{-0.13}$ 3.453.45~{}~{}3.453.45  -0.29
Table 1: Comparison of observed quark mass ratios and mixing observables at $m_{t}$ and renormalised to $10^{4}$ TeV, with those obtained by the fit to the textures, eq. (3). $J_{CP}$ is Jarlskog’s $CP$-violating invariant. The values of the fitted parameters are: $A_{0}=0.854$, $b=0.462$, $c_{d}=-0.040$, $c_{u}=0.344$ and $\lambda_{0}=0.22646$.
@@
The fit to the data renormalised at the weak scale ($\mu=m_{t}$, not shown), is excluded by its $\chi^{2}/dof$ ($\simeq 100/3$). However, the lack of observations of new physics at around this scale makes it unlikely anyway that a MM texture applies at this scale. Several of the fitted observables evolve with renormalisation scale, and the rate and direction of this evolution in the SM are well-known smRenorm1 smRenorm2 smRenorm3 ckmRenorm1 ckmRenorm2 . It is thus possible that our texture might more naturally apply at some higher energy scale, so we have scanned the SM evolution of the relevant observables up to the GUT scale, to investigate this possibility. The best fit is at a scale $\mu\sim 10^{4}$ TeV with a $\chi^{2}/dof$ [^8] While the scale $\mu$ is not a parameter of the texture, we include it in this count of fit parameters. $\simeq 1.0/2$ and a ±$\pm 1$ standard deviation width of $\sim$ a factor 3 on either side, giving the range $\mu\sim(0.3\rightarrow 3)\times 10^{4}$ TeV.
@@
The full fit results for the preferred solution are shown in table 1 where the second column shows the input data, the third gives the same inputs renormalised to the best fit scale, $\mu\sim 10^{4}$ TeV, the fourth gives the fitted values of the observables, while the fifth gives the individual pulls compared with the input observables at the best fit scale. The $\chi^{2}$ is calculated summing over those parameters above the dashed line for all observables renormalised to the best fit scale. Observables shown below the total $\chi^{2}$ line are not independent, and are not used in the fit, but are given for completeness, since the fitted values may be tested more precisely than at present by future measurements.
@@
As is usual, the signs of the masses are unphysical, so that different signs of the fitted quark mass ratios must be allowed for in the fit. Only two of the possible sign permutations have viable $\chi^{2}$ values, namely those with either sign of $m_{u}/m_{c}$, with $m_{d}/m_{s}$ negative and the other mass ratios positive. These two solutions are mutually unresolved, being almost degenerate in $\chi^{2}$, and in the values of all fit parameters except $c_{u}$, which controls $m_{u}/m_{c}$. Thus to avoid repetition, we have reported only the fit for the smaller value of $c_{u}$ (with $m_{u}/m_{c}$ negative). The fits for other permutations of mass-signs are significantly disfavoured by their $\chi^{2}$ values.
@@ 5 Constraints
where eq. (20) is a scale-dependent prediction and eqs. (21)-(22) are scale-independent. The numerical results, given for comparison above, are from table 1 and are given at the best fit scale, $\mu\sim 10^{4}$ TeV. The analytic NLO formulae are given in the Appendix, eqs. (29)-(32), and can be combined to provide a better match to the fitted values quoted above. Finally, we note the suggestive similarity between eq. (20) and the phenomenologically successful prediction for the neutrino mass ratio: $\frac{m_{2}}{m_{3}}=\tan^{2}\!{\tfrac{\pi}{8}}$ in krishnan1 .
@@ 6 Symmetries of the mass matrices
The texture, eqs. (3)-(5), and its fit, table 1, are the main results of this paper. Having established that a viable fit to the data is possible for this texture, we identify the two invariance properties of the MMs which realise the complex constant, $\bm{\lambda_{u}}/\bm{\lambda_{d}}$, defined in eq. (5).
The ratio $\bm{\lambda_{u}}/\bm{\lambda_{d}}$ involves a parameter from each MM, so that a regularity therein cannot be reflected in either matrix alone. The UT is significant in providing a diagrammatic representation of a relationship between the MMs and we will find it particularly useful in illustrating the symmetries inherent in our MM-pair. We first recall that a $CP$ transformation is effected by a simple complex conjugation: here, simply $\bm{z_{0}}\rightarrow\bm{z_{0}}^{*},~{}\bm{\overline{z}_{0}}\rightarrow\bm{\overline{z}_{0}}^{*}$. The Jarlskog invariant, $J_{CP}$, flips sign, the UT is reflected in the real axis, and other observables are preserved. Naturally, a very poor fit to the data results, due to the resulting wrong sign of $\overline{\eta}$. Less obviously, with our textures, under either $\bm{z_{0}}\rightarrow-\bm{z_{0}}$ or $\bm{\overline{z}_{0}}\rightarrow-\bm{\overline{z}_{0}}$ (but not both), we also obtain a sign-flip of $J_{CP}$ with no other consequences. Thus the simultaneous application of both $\bm{z_{0}}\rightarrow-\bm{z_{0}}$ and a $CP$ transformation cancel each other out, making the combined transformation a symmetry of our MM-pair. Analogously, simultaneous application of $\bm{\overline{z}_{0}}\rightarrow-\bm{\overline{z}_{0}}$ and $CP$ also leave the observables invariant and are another manifestation of their symmetry.
@@
where the LO-UT is again shaded yellow, its complex-conjugate green. We recall that simultaneous phase changes of the off-diagonal elements of the MMs are unobservable and correspond to rotations of the unitarity triangle in the complex plane JarlsRot1 JarlsRot2 JarlsRot3 JarlsRot4 JarlsRot5 . In the figure, we can see that a flip in the sign of $\bm{z_{0}}$ alone results in a new LO-UT (unshaded) which is congruent to the $CP$-transformed triangle, being rotatable into it. The example with the triangle formed using $\bm{z_{0}}$ and $-\bm{\overline{z}_{0}}$ behaves similarly. Figure 2 thus makes manifest the symmetry under simultaneous $CP$ transformation and single sign flip of either $\bm{\lambda_{q}}$ (but not both).
We emphasise that this symmetry is a direct result of the $\tfrac{\pi}{2}$ phase difference between $\bm{z_{0}}$ and $\bm{\overline{z}_{0}}$. It is clear from figure 2 that if $\bm{z_{0}}$ and $\bm{\overline{z}_{0}}$ had any other phase difference, then forming a triangle using $-\bm{z_{0}}$ and $\bm{\overline{z}_{0}}$ or ($\bm{z_{0}}$ and $-\bm{\overline{z}_{0}}$) would result in a triangle which was not congruent to the $CP$-transformed one, indicating that then, the symmetry is absent. One could reverse the argument and say that requiring such an invariance is sufficient to ensure that the two $\bm{\lambda_{q}}$ have a relative phase of $\tfrac{\pi}{2}$. The situation here is reminiscent of that in many models of lepton mixing wherein invariance under a combined transformation involving $CP$ and a second involution (e.g. $\mu-\tau$ exchange in mutauSymm ) results in a leptonic $CP$-phase close to $\tfrac{\pi}{2}$.
The above considerations extend to the case of the UT resulting from full diagonalisation, despite its being slightly perturbed relative to the LO one illustrated. This is verified using Jarlskog’s $CP$-violating determinant JarlskogDet1 JarlskogDet2 , which, like the UT, links both MMs:

### v4

@@ 1 Introduction
There has been a long history of textures proposed for the quark mass/Yukawa matrices (MMs), both as stand-alone ansatze, and/or as the result of (or motivation for) models which can constrain them. The traditional guiding principle for such textures has been to reduce the number of free parameters (a priori 36) to a number smaller than the ten observables, in order to find predictive constraints among the measurable quantities, especially relationships between the quark masses and mixings fritzsch1 ; fritzsch2 ; fritzsch3 ; texZero , which each manifest marked hierarchical structure. Methods have included using symmetries to enforce texture-zeroes zeroesSymms , and/or mechanisms to produce a hierarchy among the matrix elements fnMech . Recently, ideas have turned to the realm of modular forms modular0 ; modular1 , where again approaches have been found to enforce zeroes modularZeroes and/or hierarchies modularHierarchies . With the continuing advance of experimental determinations, many ansatze, e.g. fritzsch2 ; fritzsch3 , are now excluded by data, see e.g. bb . We introduce here a new MM texture based purely on the CKM and quark mass phenomenology, which is modestly predictive, fits the data, and has some novel and potentially interesting features.
The closeness of the (beauty-down) Unitarity Triangle (UT) UT1 ; UT2 to a right-angled triangle is striking masinaAndSavoy ; rightUT ; AKMS , and it is also notable that its three internal angles pdg , [^1] We use the measured value for $\beta$ and the CKM fit values pdg for $\alpha$ and $\gamma$, since these give the smallest uncertainties.
@@ 2 The texture
The proposed new texture is common to both the up and down quark MMs, constraining them relative to those of the SM. Without loss of generality, we take the MMs to be Hermitian. We adopt the popular texture zero in the $13$ and $31$ elements fritzsch2 ; fritzsch3 ; texZero , so that the smallest mixing elements result from the non-commutation of the (small) diagonalising transformations in the $12$ and $23$ subspaces. In order both to introduce viable hierarchies among the MMs’ elements, and to include $CP$ violation, we adopt two small, complex expansion parameters, [^2] Throughout this paper, where ambiguity can arise, bold font is used to denote complex quantities, and the corresponding light font symbols, their moduli. Complex conjugation is indicated by the conventional symbol. $\bm{\lambda_{u}}$ and $\bm{\lambda_{d}}$, powers of which are inserted into the up and down quark MMs respectively, each playing a role similar to Wolfenstein’s $\lambda$ parameter in the CKM matrix wolfenstein , but distinct for each of the two MMs. Thus, we express both MMs as follows:
@@
The overall normalisations of the $\bm{\lambda_{q}}~{}(q=u,d)$ are free to vary, but we will show later that their complex sum has magnitude close to the value of Wolfenstein’s expansion parameter:
@@
The constraint, eq. (5), is the most significant innovation in this paper. Specifically, in our texture, $(\bm{\lambda_{u}}/\bm{\lambda_{d}})^{*}$ is identifiable in leading approximation with the complex ratio of the two sloped sides of the UT as we show below. Thus $\arg{(-\bm{\lambda_{u}}/\bm{\lambda_{d}})}$ corresponds to the UT angle $\alpha$, the constrained value ensuring $\alpha\simeq\tfrac{\pi}{2}$; the ratio $|\bm{\lambda_{u}}/\bm{\lambda_{d}}|$ then corresponds similarly to $\tan{\beta}$ (while simultaneously determining the asymmetry between the mass hierarchies in the two charge sectors). Thus eq. (5) applied in our texture, eq. (3), yields eq. (2). Moreover, the precise values for the ratio asserted by eq. (5) imply and are implied by two novel symmetries of the $M_{q}^{0}$, as we will also show.
The MM texture, eq. (3), uses five free parameters, $A_{0}$, $b$, $c_{u}$, $c_{d}$ and $\lambda_{0}$ to fit the three quark mixing angles, the $CP$ phase, and the four quark mass ratios, making it modestly predictive. A more general parameterisation would allow the $(A_{0},b)$ parameters to be different for $q=u,d$, but we find that it is possible to fit all the quark mass and mixing observables with the parameter-pair $(A_{0},b)$, common to both mass matrices. Thus a weak isospin (up/down) reflection symmetry is respected by the real parameters ($A_{0}$, $b$), and broken only by the $\bm{\lambda_{q}}$, the pair of parameters $c_{q}$ which govern $m_{u}$ and $m_{d}$, and by the matrix normalisations, $n_{q}$ in eq. (3).
@@ 3 Leading-order solution
where $\lambda$, $\rho$, $\eta$ and $A$ are the Wolfenstein parameters wolfenstein (and eq. (13) explains our choice of symbol $A_{0}$ in eq. (3)). This confirms that at leading order, $\bm{\lambda_{u}}^{*}$ and $\bm{\lambda_{d}}^{*}$ are each proportional to one of the “sloped” sides of the UT (since, in the PDG phase convention, $V_{ud}\simeq V_{tb}\simeq 1$). We can thus immediately justify the claims made below eq. (5):
@@
$\Rightarrow\arg\left({-\frac{\bm{\lambda_{u}}}{\bm{\lambda_{d}}}}\right)$ $=\arg\left({-\frac{V_{td}}{V_{ub}^{*}}}\right)\simeq\alpha,~{}~{}{\rm and}~{}~{}\left|\frac{\bm{\lambda_{u}}}{\bm{\lambda_{d}}}\right|\simeq\tan{\beta},$ (15)
where the definition of $\alpha$ is from the PDG pdg , and the last near-equality is a consequence of the near right-angle in the resulting UT.
Still in the PDG phase convention where $\bm{\lambda_{d}}+\bm{\lambda_{u}}$ is real (and equals $\lambda_{0}$), it is useful to define the complex constants: [^5] The “bar” in $\bm{\overline{z}_{0}}$ is simply to identify this variable as the one-complement of $\bm{z_{0}}$ and should not be confused with the bar appearing above the “reduced” Wolfenstein parameters $\overline{\rho}$ and $\overline{\eta}$.
@@
The apex of the standard UT is simply given by pdg ; burasetal :
@@ 4 Numerical fit to the data
For the fit to the data, a full diagonalisation is performed at high numerical precision. The CKM observables are taken from the latest PDG summary pdg . Significant recent progress in quark mass determinations using lattice QCD flag has led to dramatic reductions in the uncertainties on the quark mass ratios (which we take as experimental inputs [^7] Lattice quark mass determinations are calibrated to experimentally-determined meson masses.). We follow the approach of bb , updating the inputs with the latest lattice values compiled in the review on quark masses pdg . These inputs are quoted in the second column of table 1 renormalised to the scale $\mu=m_{t}=173.1$ GeV in the $\overline{\rm MS}$ scheme. A $\chi^{2}$ fit to the data is performed, varying the five free parameters of the MMs to find their preferred values.
@@
$|m_{u}/m_{c}|\,(\times 10^{3})$ $2.00\pm 0.05$ Unchanged 2.002.002.00~{}~{}~{}~{}~{}~{}2.00 0.000.00~{}~{}0.000.00
$|m_{d}/m_{s}|\,(\times 10^{2})$ $4.97\pm 0.06$ 4.974.974.97~{}~{}~{}~{}~{}~{}4.97 0.030.03~{}~{}0.030.03
$m_{c}/m_{t}\,(\times 10^{3})$ $3.67\pm 0.04$ $3.46\pm 0.03$ 3.463.463.46~{}~{}~{}~{}~{}~{}3.46 $-0.01$
$m_{s}/m_{b}\,(\times 10^{2})$ $1.854\pm 0.007$ $1.968\pm 0.008$  1.9681.968~{}\,1.968~{}~{}~{}~{}~{}~{}1.968 0.020.02~{}~{}0.020.02
$\lambda$ $0.22501\pm 0.00068$ Unchanged 0.224990.22499~{}~{}~{}~{}0.22499~{}~{}~{}~{}~{}~{}0.22499 $-0.03$
$A$ $0.826^{+0.016}_{-0.015}$ $0.877^{+0.017}_{-0.016}$ 0.8760.876~{}0.876~{}~{}~{}~{}~{}~{}0.876 $-0.09$
$\overline{\rho}$ $0.1591\pm 0.0094$ Unchanged 0.15190.1519~{}~{}~{}0.1519~{}~{}~{}~{}~{}~{}0.1519 $-0.76$
$\overline{\eta}$ $0.3523^{+0.0073}_{-0.0071}$ 0.34770.3477~{}~{}~{}0.3477~{}~{}~{}~{}~{}~{}0.3477 $-0.65$
@@
$\beta\,(^{\circ})$ $22.6\pm 0.4$  ${}\,22.3\pm 0.1$ $-0.75$
$\gamma\,(^{\circ})$ $65.7\pm 1.3$  ${}\,66.4\pm 0.1$  0.540.54~{}~{}\,0.540.54
$\,\beta_{s}\,(^{\circ})$ $1.15\pm 0.46$ ${}~{}~{}1.07\pm 0.01$ $\,-0.17$
Table 1: Comparison of observed quark mass ratios and mixing observables at $m_{t}$ and renormalised to $10^{4}$ TeV, with those obtained by the fit to the textures, eq. (3). The predictions shown for the UT angles in the penultimate column apply also at the weak scale, since they have negligible evolution with renormalisation scale. $\beta_{s}$ is determined from $CP$-violation in $B^{0}_{s}$ decays pdg . Pulls are calculated as the difference between the central fitted value and the renormalised input values, as a fraction of the renormalised experimental errors. The values of the fitted parameters are: $A_{0}=0.854\pm 0.013$, $b=0.462\pm 0.001$, $c_{d}=-0.040\pm 0.006$, $c_{u}=0.344\pm 0.003$ and $\lambda_{0}=0.22646\pm 0.00034$.
@@
The fit to the data renormalised at the weak scale ($\mu=m_{t}$, not shown), is excluded by its $\chi^{2}/dof$ ($\simeq 100/3$). However, the lack of observations of new physics at around this scale makes it unlikely anyway that a MM texture applies at this scale. Several of the fitted observables ($A$, $m_{c}/m_{t}$ and $m_{s}/m_{b}$) evolve with renormalisation scale, and the rate and direction of this evolution in the SM are well-known smRenorm1 ; smRenorm2 ; smRenorm3 ; ckmRenorm1 ; ckmRenorm2 . It is thus possible that our texture might more naturally apply at some higher energy scale, so we have scanned the SM evolution of the relevant observables up to the GUT scale, to investigate this possibility. The best fit is at a scale $\mu\sim 10^{4}$ TeV with a $\chi^{2}/dof$ [^8] While the scale $\mu$ is not a parameter of the texture, we include it in this count of fit parameters. $\simeq 1.0/2$ and a width of $\sim$ a factor 3 on either side, giving the range $\mu\sim(0.3\rightarrow 3)\times 10^{4}$ TeV.
@@
The full fit results for the preferred solution are shown in table 1 where the second column shows the input data, the third gives the same inputs renormalised to the best fit scale, $\mu\sim 10^{4}$ TeV, the fourth gives the fitted values of the observables, while the fifth gives the individual pulls compared with the input observables at the best fit scale. The $\chi^{2}$ is calculated summing over those parameters above the dashed line for all observables renormalised to the best fit scale. The best fit parameter values and their errors are given in the table caption. Observables shown below the total $\chi^{2}$ line are not independent, and are not used in the fit, but are given for completeness as scale-independent predictions. Their values may be tested more precisely than at present by future measurements.
@@
As is usual, the signs of the masses are unphysical, so that different signs of the fitted mass ratios must be allowed for in the fit. Only two of the possible sign permutations have viable $\chi^{2}$ values, i.e. those with either sign of $m_{u}/m_{c}$, with $m_{d}/m_{s}$ negative and the other mass ratios positive. These two solutions are mutually unresolved, being almost degenerate in $\chi^{2}$, and in the values of all fit parameters except $c_{u}$, which controls $m_{u}/m_{c}$. To avoid repetition, we have reported only the fit for the smaller value of $c_{u}$ (with $m_{u}/m_{c}$ negative). The fits for other mass-sign permutations are strongly disfavoured by their $\chi^{2}$ values.
@@ 5 Constraints
where eq. (20) is a scale-dependent prediction and eqs. (21)-(22) are scale-independent. The numerical results, given for comparison above, are from table 1 and are given at the best fit scale, $\mu\sim 10^{4}$ TeV. The analytic NLO formulae are given in the Appendix, eqs. (29)-(32), and can be combined to provide a better match to the fitted values quoted above. Finally, we note the suggestive similarity between eq. (20) and the phenomenologically-successful prediction for the neutrino mass ratio: $\frac{m_{2}}{m_{3}}=\tan^{2}\!{\tfrac{\pi}{8}}$ in krishnan1 .
@@ 6 Symmetries of the mass matrices
The texture, eqs. (3) and (5), and its fit, table 1, are the main results of this paper. Having established that a viable fit to the data is possible for this texture, we identify the two invariance properties of the MMs which realise the complex constant, $\bm{\lambda_{u}}/\bm{\lambda_{d}}$, defined in eq. (5).
The ratio $\bm{\lambda_{u}}/\bm{\lambda_{d}}$ involves a parameter from each MM, so that a regularity therein cannot be reflected in either matrix alone. The UT is significant in providing a diagrammatic representation of a relationship between the MMs and we will find it particularly useful in illustrating the symmetries inherent in our MM-pair. We first recall that a $CP$ transformation is effected by a simple complex conjugation of the MMs: here, simply $\bm{z_{0}}\rightarrow\bm{z_{0}}^{*},~{}\bm{\overline{z}_{0}}\rightarrow\bm{\overline{z}_{0}}^{*}$. In consequence, the Jarlskog invariant, $J_{CP}$, flips sign, the UT is reflected in the real axis, and other observables are preserved. Naturally, a very poor fit to the data results, due to the resulting wrong sign of $CP$-violating observables. Less obviously, with our textures, under either $\bm{z_{0}}\rightarrow-\bm{z_{0}}$ or $\bm{\overline{z}_{0}}\rightarrow-\bm{\overline{z}_{0}}$ (but not both), we also obtain a sign-flip of $J_{CP}$, and effectively (see below) a reflection of the UT in the real axis. Thus the simultaneous application of both $\bm{z_{0}}\rightarrow-\bm{z_{0}}$ and a $CP$ transformation cancel each other out, making the combined transformation a symmetry of our MM-pair. Analogously, the simultaneous application of $\bm{\overline{z}_{0}}\rightarrow-\bm{\overline{z}_{0}}$ together with $CP$ likewise leaves the observables invariant, an alternative manifestation of the same symmetry.
@@
where the LO-UT is again shaded yellow, with its complex-conjugate (i.e. its $CP$-transform) green. We recall that simultaneous phase changes of the off-diagonal elements of the MMs are unobservable and correspond to rotations of the unitarity triangle in the complex plane JarlsRot1 ; JarlsRot2 ; JarlsRot3 ; JarlsRot4 ; JarlsRot5 . In the figure, we can see that a flip in the sign of $\bm{z_{0}}$ alone results in a new LO-UT (unshaded) which is congruent to the $CP$-transformed triangle, being rotatable into it. The example with the triangle formed using $\bm{z_{0}}$ and $-\bm{\overline{z}_{0}}$ behaves similarly. Figure 2 thus makes manifest the symmetry under simultaneous $CP$ transformation and single sign flip of either $\bm{\lambda_{q}}$ (but not both).
We emphasise that this symmetry is a direct result of the $\tfrac{\pi}{2}$ phase difference between $\bm{z_{0}}$ and $\bm{\overline{z}_{0}}$. It is clear from figure 2 that if $\bm{z_{0}}$ and $\bm{\overline{z}_{0}}$ had any other phase difference, then forming a triangle using $-\bm{z_{0}}$ and $\bm{\overline{z}_{0}}$ or ($\bm{z_{0}}$ and $-\bm{\overline{z}_{0}}$) would result in a triangle which was not congruent to the $CP$-transformed one, indicating that then, the symmetry is absent. One could reverse the argument and say that requiring such an invariance in the MMs is sufficient to ensure that the two $\bm{\lambda_{q}}$ have a relative phase of $\tfrac{\pi}{2}$. The situation here is reminiscent of that in many models of lepton mixing wherein invariance under a combined transformation involving $CP$ and a second involution (e.g. $\mu-\tau$ exchange in mutauSymm ) results in a leptonic $CP$-phase close to $\tfrac{\pi}{2}$.
The above considerations extend to the case of the UT resulting from full diagonalisation, despite its being slightly perturbed relative to the LO-UT illustrated. This is verified using Jarlskog’s $CP$-violating determinant JarlskogDet1 ; JarlskogDet2 , which, like the UT, links both MMs: